OpenStudy (anonymous):

what is the equation of the line that passes through the points (0 1) and (6 -2)

OpenStudy (sidsiddhartha):

let (x1,y1)=(0,1) and (x2,y2)=(6,-2) so the equation of the line will be (x-x1)/(x1-x2)=(y-y1)/(y1-y2)